OpenStudy (anonymous):

Solve. 8y – 24 > –16 A. y > –5 B. y > –1 C. y > 1 D. y > 5

OpenStudy (anonymous):

@AriPotta

sammixboo (sammixboo):

First add 24 on both sides of the equation 8y - 24 + 24 = -16 + 24 8y = -16 + 24 Can you tell me what is -16 + 24 is?

OpenStudy (anonymous):

-40

sammixboo (sammixboo):

Nope, try again :)

OpenStudy (anonymous):

8

sammixboo (sammixboo):

Correct! So now we have 8y = 8, now we must divide 8 on both sides of the equation 8y/8 = 8/8 y = 8/8 What is 8/8?

OpenStudy (anonymous):

1

sammixboo (sammixboo):

So y > 1